# ACT inflation and its influence on reheating era in Einstein-Gauss-Bonnet gravity

### v1

@@ I Introduction
Currently we are living in a cosmological era which has witnessed a giant leap of understanding on early universe phenomenon, primarily from Cosmic Microwave Background (CMB) CMB-S4:2016ple ; Planck:2018jri . The recent Atacama Cosmology Telescope (ACT) data, combined with the Planck 2018 and the Baryon Accoustoc Oscillation (BAO) from DESI, constrain the inflationary models in a way that the curvature perturbation should be less red tilted ACT:2025fju ; ACT:2025tim . In particular, the Data Release 6 (DR6) from the ACT, combined with the Planck 2018 and BAO (collectively it can be denoted by ACT-DR6+Planck18+BAO), refines the constraint on inflationary observables (like the spectral index of primordial curvature perturbation and tensor-to-scalar ratio) as: $n_{s}=0.9743\pm 0.0034$ and $r<0.038$ respectively ACT:2025fju ; ACT:2025tim , indicating a $2\sigma$ deviation from the only-Planck 2018 measurements Planck:2018jri . Such a refinement on $n_{s}$ and $r$ from ACT data disfavors the models which previously lie at lower boundary value of $n_{s}$, however allows such models which produce higher value of $n_{s}$ such as $\alpha$-attractor inflationary model. Based on these argument, a variety of inflationary models (which are viable from only-Planck 2018 result) are revisited from the angle of the recent ACT data Kallosh:2025rni ; Aoki:2025wld ; Odintsov:2025wai ; Haque:2025uga ; Dioguardi:2025vci ; Salvio:2025izr ; Antoniadis:2025pfa ; Kim:2025dyi ; Gao:2025onc ; Drees:2025ngb .
@@
The scalar-Einstein-GB theory is motivated due to its rich cosmological consequences starting from early inflation to the consistency with the DESI data regarding the late dark energy era Nojiri:2005vv ; Li:2007jm ; Odintsov:2020sqy ; Cognola:2006eg ; Odintsov:2023lbb ; Odintsov:2018zhw ; Chakraborty:2018scm ; Kanti:2015pda ; Elizalde:2020zcb ; Pozdeeva:2021iwc ; Granda:2019jqy ; Nojiri:2022xdo ; Oikonomou:2024jqv ; Nojiri:2023mbo ; Odintsov:2025kyw . We should mention that the Gauss-Bonnet inflation in respect to the ACT data has been recently studied in Odintsov:2025wai , however in quite different contexts. Note that, in our present analysis, we include the post-inflationary reheating dynamics to examine how the ACT data influence the reheating phenomenology as well in Gauss-Bonnet cosmology, which makes the present scenario essentially different from earlier one Odintsov:2025wai where only the inflationary scenario is taken into account. Moreover Odintsov:2025wai put an extra condition on the Gauss-Bonnet coupling function to make the model compatible with GW170817 event which states that the speed of gravitational waves is equal to the speed of light. However in our current paper, the model is compatible with GW170817 event, but due to the presence of reheating stage rather than incorporating any new condition on the Gauss-Bonnet coupling. Actually the speed of the gravitational wave during inflation in scalar-Einstein-GB theory is different than the speed of light (which is unity in natural units). However at the late time, in particular after the reheating phase, the scalar field energy density transforms to relativistic particles leading to the radiation dominated era with a suitable reheating temperature. In this regard, if we consider that the radiation energy does not exhibit any non-minimal coupling with the GB term, then the higher curvature GB term becomes a surface term in the four dimensional gravitational action after the reheating phase. As a consequence, the scalar-Einstein-GB gravity effectively reduces to the standard Einstein gravity with a perfect fluid (radiation) after the reheating phase, that leads to unit gravitational speed — which is indeed consistent with GW170817 event.

### v2

@@ I Introduction
Currently we are living in a cosmological era which has witnessed a giant leap of understanding on early universe phenomenon, primarily from Cosmic Microwave Background (CMB) CMB-S4:2016ple ; Planck:2018jri . The recent Atacama Cosmology Telescope (ACT) data, combined with the Planck 2018 and the Baryon Accoustoc Oscillation (BAO) from DESI, constrain the inflationary models in a way that the curvature perturbation should be less red tilted ACT:2025fju ; ACT:2025tim . In particular, the Data Release 6 (DR6) from the ACT, combined with the Planck 2018 and BAO (collectively it can be denoted by ACT-DR6+Planck18+BAO), refines the constraint on inflationary observables (like the spectral index of primordial curvature perturbation and tensor-to-scalar ratio) as: $n_{s}=0.9743\pm 0.0034$ and $r<0.038$ respectively ACT:2025fju ; ACT:2025tim , indicating a $2\sigma$ deviation from the only-Planck 2018 measurements Planck:2018jri . Such a refinement on $n_{s}$ and $r$ from ACT data disfavors the models which previously lie at lower boundary value of $n_{s}$, however allows such models which produce higher value of $n_{s}$ such as $\alpha$-attractor inflationary model. Based on these argument, a variety of inflationary models (which are viable from only-Planck 2018 result) are revisited from the angle of the recent ACT data Kallosh:2025rni ; Aoki:2025wld ; Odintsov:2025wai ; Haque:2025uga ; Dioguardi:2025vci ; Salvio:2025izr ; Antoniadis:2025pfa ; Kim:2025dyi ; Gao:2025onc ; Drees:2025ngb ; Maity:2025czp ; Mondal:2025kur ; Peng:2025bws ; Liu:2025qca .
@@
The scalar-Einstein-GB theory is motivated due to its rich cosmological consequences starting from early inflation to the consistency with the DESI data regarding the late dark energy era Nojiri:2005vv ; Satoh:2008ck ; Satoh:2007gn ; Li:2007jm ; Odintsov:2020sqy ; Cognola:2006eg ; Odintsov:2023lbb ; Odintsov:2018zhw ; Chakraborty:2018scm ; Kanti:2015pda ; Elizalde:2020zcb ; Pozdeeva:2021iwc ; Granda:2019jqy ; Nojiri:2022xdo ; Oikonomou:2024jqv ; Nojiri:2023mbo ; Odintsov:2025kyw ; Hussain:2025vbo . We should mention that the Gauss-Bonnet inflation in respect to the ACT data has been recently studied in Odintsov:2025wai , however in quite different contexts. Note that, in our present analysis, we include the post-inflationary reheating dynamics to examine how the ACT data influence the reheating phenomenology as well in Gauss-Bonnet cosmology, which makes the present scenario essentially different from earlier one Odintsov:2025wai where only the inflationary scenario is taken into account. Moreover Odintsov:2025wai put an extra condition on the Gauss-Bonnet coupling function to make the model compatible with GW170817 event which states that the speed of gravitational waves is equal to the speed of light. However in our current paper, the model is compatible with GW170817 event, but due to the presence of reheating stage rather than incorporating any new condition on the Gauss-Bonnet coupling. Actually the speed of the gravitational wave during inflation in scalar-Einstein-GB theory is different than the speed of light (which is unity in natural units). However at the late time, in particular after the reheating phase, the scalar field energy density transforms to relativistic particles leading to the radiation dominated era with a suitable reheating temperature. In this regard, if we consider that the radiation energy does not exhibit any non-minimal coupling with the GB term, then the higher curvature GB term becomes a surface term in the four dimensional gravitational action after the reheating phase. As a consequence, the scalar-Einstein-GB gravity effectively reduces to the standard Einstein gravity with a perfect fluid (radiation) after the reheating phase, that leads to unit gravitational speed — which is indeed consistent with GW170817 event.
@@ III Inflationary era
On a different perspective, here it may be mentioned that the author of Pozdeeva:2020shl considered a different set of scalar potential and GB coupling function in order to examine the attractor behaviour along with observational viability (during inflation) of scalar-Einstein-GB theory. In particular, Pozdeeva:2020shl used the following sets of $V(\phi)$ and $\xi(\phi)$:
$V(\phi)\sim\mathrm{exp}\left[-2\mathrm{e}^{-p\phi}\right]~~~~~~~\mathrm{and}~~~~~~\xi(\phi)\sim\mathrm{exp}\left[2\mathrm{e}^{-p\phi}\right]\,$, (29)
where $p$ is a constant. For such $V(\phi)$ (and $\xi(\phi)$), the inflationary parameters (like $n_{\mathrm{s}}$ and $r$) in the large field regime depends on the factor $\mathrm{e}^{-p\phi}$ Pozdeeva:2020shl . As a result, it turns out that with the condition: $\mathrm{e}^{-p\phi}\ll 1$, the above scalar-Einstein-GB model (in Eq. (29)) becomes a viable one.